If the population consists of numbers 1, 2, 3, 4, 5, 6 and 7, find the mean of the samples of size 2 without replacement. Construct a sampling distribution.


1
Expert's answer
2021-06-08T11:01:51-0400

N=7

n=2

Number of samples "= \frac{N!}{n!(N-n)!}"

"= \frac{7!}{2!(7-2)!}=\frac{42}{2}=21"